# Bridging the Gap: Towards an Expanded Toolkit for ML-Supported Decision-Making in the Public Sector

### v2

@@ 2.2 Label Bias
Obtaining accurate ground truth data for the prediction target in real-world settings is rarely simple. Frequently, the true target is not directly measurable (Coston et al., , 2023; Guerdan et al., 2023b, ; Barocas et al., , 2019), as is the case with health outcomes or crime rates. This often encourages the use of proxy variables that are more easily available, such as hospitalization records and arrest rates. Similarly, it may take a considerable amount of time before the outcome of interest can be observed; predicting the 10-year risk of cardiovascular events takes at least a decade. Such lag may require the use of short-term outcomes as proxies (Athey et al., 2019a, ).
However, using proxy variables can introduce bias into a model. Proxy variables often capture institutional responses rather than the true underlying outcome of interest. For example, using ICU hospitalization as a proxy for COVID-19 severity is an imperfect measure, as ICU admission will depend on other factors like bed availability and other admission criteria. This can be especially problematic if the relationship between proxy and true target varies by protected attributes, such as race and gender (Guerdan et al., 2023b, ; Passi and Barocas, , 2019). Obermeyer et al., (2019) demonstrate that using expected healthcare cost as a proxy for health needs in predictive algorithms can lead to significantly underestimating the risk score of Black patients. This is because Black patients with similar health needs generate fewer medical expenditures compared to white patients. Similar examples can be found in various application contexts, such as judicial bail prediction (Fogliato et al., , 2020) and lending algorithms (Mitchell et al., , 2021).
@@ 2.3 Past Decision-Making
When developing an ADM system, we often encounter scenarios in which the available training has been influenced by past decision-making (Coston et al., , 2020). For example, a model predicting the risk of jobseekers becoming long-term unemployed with the aim of allocating future support programs needs to account for how such programs were distributed in the past. Otherwise the model will likely underestimate the risk of unemployment for individuals that used to receive prioritized support after the decision-making process is altered through the deployment of the model (Lenert et al., , 2019). The predictions of such a model would lead to misleading recommendations, since its predictions are valid only under the assumption that the decision-making policies remain unchanged or that the interventions were largely ineffective in the past (Dickerman and Hernán, , 2020).
@@
The canonical way of dealing with such bias are randomized controlled trials (RCTs) (Caron et al., , 2022). However, in many scenarios it will be impossible to conduct a RCT due to resource constraints or ethical limitations (Caron et al., , 2022). When the efficacy of the interventions is well-established, a randomized study may be hard to justify, as in the case of criminal justice (Lakkaraju et al., , 2017) or child abuse prevention (Vaithianathan et al., , 2021).
Alternatively, causal outcomes may be estimated from observational data. However, this requires the assumption that relevant confounding variables have been observed, allowing for the disentanglement of past intervention assignment and outcomes. However, some variables may remain elusive (Rambachan et al., , 2022; Lakkaraju et al., , 2017), such as the impressions gained by decision-makers from in-person interactions. In situations in which it is difficult to guarantee no unmeasured confounding variables, there still may be ways to estimate the outcome of interest. For instance, Chen et al., (2023) and Lakkaraju et al., (2017) utilize data from multiple randomly assigned human decision-makers who were randomly assigned to cases to enable estimation.
@@ 2.4 Competing Objectives and Constraints
Clearly specifying the optimization targets of an ADM system is a delicate process that carries the risk of distorting the originally intended goals (Levy et al., , 2021). This risk is especially pronounced when some policy goals are easier to formalize than others, prompting the oversimplification of complex issues through an algorithmic lens (Levy et al., , 2021). Stakeholders’ preference for cost-effective, straightforward solutions and easily measurable prediction targets may exacerbate this problem (Barocas et al., , 2019). Nevertheless, decisions must be made, and the growing use of ML in the public sector will likely require new dialogues among stakeholders, while also providing an opportunity to make the weighting and tradeoffs between policy objectives more explicit and transparent than in the past (Coyle and Weller, , 2020; Levy et al., , 2021).
@@ 2.5 Human-in-the-Loop
In many critical application contexts, the goal is not to entirely replace humans with ADM systems, but rather to aid human decision-makers in their tasks (Enarsson et al., , 2022). Human oversight becomes particularly important in scenarios where the ML system alone can not fulfill all the necessary system requirements (Doshi-Velez and Kim, , 2017), as determined by the underlying policy goals. Human quality control can be essential to ensure both system reliability and safety, especially when quantifying and anticipating model uncertainties proves challenging (Gruber et al., , 2023). Additionally, as the definition of an ADM system’s objective can often be difficult, human input can aid in balancing complex and hard to formalize objectives.
These issues are often reflected in legal and regulatory restrictions. For example, the current EU proposal for an AI legal framework stresses the importance of human oversight for high-risk systems in Article 14 (EU Commission, , 2021). In such scenarios, humans maintain the final say in determining the right course of action, necessitating considerations of how they will interpret model outputs and incorporate them into their judgment. Generally, this requires shifting the focus from estimating the most accurate model to evaluating the consequences of providing a human decision-maker with a particular recommendation (Fernández-Loría and Provost, 2022b, ; Vodrahalli et al., , 2022).
@@
For the interaction between human decision-makers and ML models to work, model predictions and its functioning must be comprehensible for human users (Yeomans et al., , 2019; Nourani et al., , 2019; Amarasinghe et al., , 2023). For instance, Lebovitz et al., (2022) show how opaque ML models make it more difficult for medical professional to effectively use them for diagnosis. Many methods have been proposed to make ML models interpretable and explainable, with comprehensive overviews available in (Belle and Papantonis, , 2021; Molnar, , 2022; Doshi-Velez and Kim, , 2017; Murdoch et al., , 2019; Rudin et al., , 2022). One of the reasons why these approaches vary widely is because they have very different conceptions of what constitutes an understandable explanation of the output of a model. Amarasinghe et al., (2023) establish an initial taxonomy linking public policy use cases with existing explainable ML approaches. Moving forward, they stress the need to rigorously evaluate explainable ML methods in real-world problem contexts to ensure their effectiveness in achieving real policy goals and in aiding domain experts. Research from fields such as psychology, cognitive sciences and philosophy may help in the task of creating explanations that are helpful to human users (Miller, , 2019). This requires careful investigation of various challenges, such as identifying situations where model explanations may be harmful due to information overload (Poursabzi-Sangdeh et al., , 2021), or when users may take advantage of increased transparency to exploit a system (Molnar, , 2022). Similarly, misleading explanations can be used to manipulate users and unjustifiably increase trust in a system (Lakkaraju and Bastani, , 2020).
@@
3 Expanding the ADM Toolkit: From Allocation Principle to Target Estimand
@@
Successfully constructing ML systems to inform decision-making and addressing the outlined challenges is a fundamentally difficult task that can not easily be tackled in the traditional supervised ML framework. Consequently, there have been calls to move beyond purely predictive modeling towards ML methodology that centers around the goal of decision-making (Hüllermeier, , 2021). This involves a shift in perspective from solely focusing on achieving accurate predictions to a more holistic modus operandi centered on selecting a modeling approach that can best inform the decision-making for a given policy goal. However, selecting an appropriate modeling approach first requires an understanding of which target estimand can best inform decision-making for a given project.
@@
This task is complex, as it can be challenging to clarify the intended guiding principles of an ADM system even when certain goals are already set. For example, public employment agencies often seek to allocate resources to jobseekers at higher risk of long-term unemployment. However, this objective might stem from either a belief in the effectiveness of early interventions for high-risk jobseekers or the notion that high-risk jobseekers inherently deserve more support (Desiere and Struyven, , 2021). Such ambiguity can present difficulties, complicating the choice of the appropriate target estimand, as a need-based distribution necessitates different modeling considerations than an approach focused on the most efficient allocation of interventions.
In practice, this discussion becomes even more complicated than the scope we can cover here, given that decision-making often extends beyond binary interventions at a specific point in time (Hüllermeier, , 2021). This complexity may require alternative target estimands, especially in scenarios involving multi-valued and continuous interventions that can simultaneously target multiple risk factors or occur sequentially at different time points (Lin et al., , 2021; Acharki et al., , 2023; Van Geloven et al., , 2020). Recent works have emphasized the importance of clearly defining the target estimand (Lundberg et al., , 2021). For example, Lin et al., (2021) and Van Geloven et al., (2020) review various estimands and modeling approaches for clinical prediction models. Clear definition of the target of interest is of critical importance, context-dependent and should align with the requirements of the decision-making process.
@@
In the following sections, we will explore classical risk prediction (Section 3.1), counterfactual prediction (Section 3.2) and policy learning (Section 3.3) as key instances of target estimands of an ADM system. We show in which context which estimand is particularly relevant, especially when the modeling of past decision-making is required (Section 2.3), and discuss different estimation strategies for each. We then present methodological advancements for each modeling approach with which the challenges described in the previous section can be addressed in a given estimation setting. Specifically, we focus on distribution shifts to ensure robustness across deployment environments, uncertainty quantification as a key building block for generating trustworthy predictions for humans, and multi-objective optimization to manage tradeoffs between competing objectives.
@@ 3.1.1 Distribution Shifts, Selection Bias and Label Bias
Most supervised learning models assume that training and deployment data follow the same distribution. However, in many real-world scenarios we may encounter a distribution shift between the training and deployment environment, necessitating the development of reliable models capable of handling and mitigating such differences (Duchi and Namkoong, , 2021). Training models that remain valid under distribution shift often requires assumptions about the expected type of shift (David et al., , 2010). As outlined in Section 2.1, we will predominantly focus on on covariate, label and concept shits (Moreno-Torres et al., , 2012; Quinonero-Candela et al., , 2008). Additionally, we will explore label bias as a type of distribution shift introduced through the use of proxy variables, and consider shifts in time caused by non-stationary environments. We outline central research streams below – for more in-depth introductions to transfer learning, domain adaptation and out-of-distribution generalization approaches, see Kouw and Loog, (2018) and Zhou et al., (2022).
The survey research literature is an invaluable resource for understanding and systematizing different error sources in the data collection and curation process that may surface as distribution shifts downstream. With their inherent focus on valid population inference, concepts such as under-coverage (relevant subpopulations cannot be reached with a data collection schema) and non-response (potentially selective non-participation of relevant subpopulations) extend beyond the traditional survey setting and can help to identify and systematize deficits in the training data composition in relation to the target population. Error frameworks such as the Total Survey Error (Groves and Lyberg, , 2010; Biemer, , 2010) and its adaptions to online data sources (Sen et al., , 2021) have been proposed to systematically trace errors along the data collection and processing pipeline that may accumulate in misrepresentation (and measurement error) in the eventual data available for model training.
Reflecting the growing use of alternative data sources for population inference, a parallel body of work proposes strategies for improving inference from data that was not collected using traditional sampling schemes (Cornesse et al., , 2020; Yang and Kim, , 2020). In this context, pseudo-weighting approaches (Elliott and Valliant, , 2017; Valliant and Dever, , 2011) are employed to match the potentially biased source data to some known reference distribution, which resembles methodology from the domain adaptation literature (see below) and similarly connects to concepts in causal inference (Mercer et al., , 2017). As a recent example of cross-disciplinary work in this context, Kim et al., (2022) draw on the multicalibration framework from algorithmic fairness (Hebert-Johnson et al., , 2018) to learn prediction functions that are universally adaptable to unknown deployment shifts and allow for downstream inference that is competitive with target-specific correction techniques.
@@
However, weighting methods struggle in settings with limited and complex source data, frequently resulting in high variance estimates, and depend on data being available from the target domain of interest (Fang et al., , 2020; Kouw and Loog, , 2018; Liu and Ziebart, , 2014). Distributionally robust methods offer an alternative approach by providing worst-case guarantees. They often involve minimax estimation, which seek to minimize the loss under the least favorable distribution shift (Kouw and Loog, , 2018; Subbaswamy et al., , 2022; Duchi and Namkoong, , 2021). For instance, Wen et al., (2014) focus on minimizing loss for the worst-case re-weighting of the source data within a predefined set of re-weighting functions.
@@
Various approaches have been devised to mitigate label noise and measurement error. For instance, Natarajan et al., (2013) propose an unbiased risk minimization strategy for handling class-conditional $p(\tilde{Y}|Y)$ noise. While such a simplified model of a proxy may be applicable in some settings, practitioners will likely encounter more complex scenarios, such as the measurement error depending on sensitive covariates (Obermeyer et al., , 2019). In many situations, there may be the option to access multiple proxies of the true target of interest. For example, Boeschoten et al., (2021) utilize a structural equation model to characterize the relationship between multiple proxies and the unobserved outcome to ensure fair predictions. There has also been research into how label bias interacts with other distribution shifts. For example, Dai and Brown, (2020) propose a joint framework for addressing label bias and label shift, while Yu et al., (2020) investigate the interaction between class-conditional noise and generalized target shifts.
Distribution shifts tend to occur gradually over time (Webb et al., , 2016), often triggered by the deployment of the model itself. Addressing such feedback loops and ongoing distribution shifts poses a significant challenge, likely requiring future research into the temporal dynamics of ML-informed decision-making (Pagan et al., , 2023). For example, Perdomo et al., (2020) introduce a modeling framework that incorporates the potential impact of predictions on the predicted outcome of interest. These predictions are referred to as performative, effectively leading to distribution shifts by altering the target distribution in the deployment environment over time. They develop the notion of performative optimality, ensuring that a decision rule minimizes the expected loss with regard to the future target distribution it induces. The specific choice of the loss function can align with different objectives. For instance, one may opt to optimize for a target distribution with mostly favorable outcomes instead of solely focusing on accurate predictions (Kim and Perdomo, , 2022).

### v3

@@ 2.2 Label Bias
Obtaining accurate ground truth data in real-world settings is rarely simple, as the true quantity of interest is often not directly measurable (Coston et al., , 2023; Guerdan et al., 2023b, ; Barocas et al., , 2023). While challenges in measuring outcomes are not unique to the public sector, they are particularly pronounced in the public sector, where projects often address complex social phenomena that are difficult to quantify such as health, social welfare and education. In contrast, the private sector typically evaluates outcomes using more straightforward metrics like return on investment (Wirick, , 2011). These difficulties often encourage the use of proxy variables that are more easily available, such as hospitalization records and arrest rates. Similarly, it may take a considerable amount of time before the outcome of interest can be observed; predicting the 10-year risk of cardiovascular events takes at least a decade. Such lag may require the use of short-term outcomes as proxies (Athey et al., 2019a, ).
Using proxy variables can introduce bias into a model. Proxy variables often capture institutional responses rather than the true underlying outcome of interest. For example, using ICU hospitalization as a proxy for COVID-19 severity is an imperfect measure, as ICU admission will depend on other factors like bed availability and other admission criteria. This can be especially problematic if the relationship between proxy and true target varies by protected attributes, such as race and gender (Guerdan et al., 2023b, ; Passi and Barocas, , 2019). Obermeyer et al., (2019) demonstrate that using expected healthcare cost as a proxy for health needs in predictive algorithms can lead to significantly underestimating the risk score of Black patients. This is because Black patients with similar health needs generate fewer medical expenditures compared to white patients. Similar examples can be found in various application contexts, such as judicial bail prediction (Fogliato et al., , 2020) and lending algorithms (Mitchell et al., , 2021).
@@ 2.3 Past Decision-Making
When developing an ADM system, we often encounter scenarios in which the available training data has been influenced by past decision-making (Coston et al., , 2020). For example, a model predicting the risk of job seekers becoming long-term unemployed with the aim of allocating future support programs needs to account for how such programs were distributed in the past. Otherwise the model will likely underestimate the risk of unemployment for individuals that used to receive prioritized support after the decision-making process is altered through the deployment of the model (Lenert et al., , 2019). The predictions of such a model would lead to misleading recommendations, since its predictions are valid only under the assumption that the decision-making policies remain unchanged or that the interventions were largely ineffective in the past (Dickerman and Hernán, , 2020).
@@
The canonical way of dealing with such bias are randomized controlled trials (RCTs) (Caron et al., , 2022). However, in many high-stakes public sector settings it will be impossible to conduct a RCT due to resource constraints and ethical limitations (Caron et al., , 2022). When the efficacy of the interventions is well-established, a randomized study may be hard to justify, as in the case of criminal justice (Lakkaraju et al., , 2017) or child abuse prevention (Vaithianathan et al., , 2021).
Alternatively, causal outcomes may be estimated from observational data. However, this requires the assumption that relevant confounding variables have been observed, allowing for the disentanglement of past intervention assignment and outcomes. However, some variables may remain elusive (Rambachan et al., , 2022; Lakkaraju et al., , 2017), such as the impressions gained by decision-makers from in-person interactions. In situations in which it is difficult to guarantee no unmeasured confounding variables, there still may be ways to estimate the outcome of interest. For instance, Chen et al., (2023) and Lakkaraju et al., (2017) utilize data from multiple human decision-makers who were randomly assigned to cases to enable estimation.
@@ 2.4 Competing Objectives and Constraints
Clearly specifying the optimization targets of an ADM system is a delicate process that carries the risk of distorting the originally intended goals (Levy et al., , 2021). This risk is especially pronounced when some policy goals are easier to formalize than others, prompting the oversimplification of complex issues through an algorithmic lens (Levy et al., , 2021). Stakeholders’ preference for cost-effective, straightforward solutions and easily measurable prediction targets may exacerbate this problem (Barocas et al., , 2023). Nevertheless, decisions must be made, and the growing use of ML in the public sector will likely require new dialogues among stakeholders, while also providing an opportunity to make the weighting and tradeoffs between policy objectives more explicit and transparent than in the past (Coyle and Weller, , 2020; Levy et al., , 2021).
@@ 2.5 Human-in-the-Loop
Automated systems alone often cannot meet all the criteria necessary for real-world deployment, such as ensuring reliability under unexpected conditions, transparency and accountability. This makes integrating human decision-makers with algorithmic systems a central concern in the public sector, where systems inform high-stakes decisions and need to comply with complex regulatory frameworks. Mitrou et al., (2022) highlight the need for human discretion and oversight when systems continuously learn on (biased) historical data, face competing objectives and values, or need to meet accountability obligations. These concerns are often reflected in legal frameworks like the EU’s proposal for AI regulation, which stresses the importance of human oversight for high-risk systems in Article 14 (EU Commission, , 2021). Thus, the goal is not to replace humans with ADM systems but to assist public decision-makers in their tasks (Enarsson et al., , 2022).
In such scenarios, humans must maintain the final say, making it important to consider how they interpret model outputs and integrate them into their decision-making process. This shifts the focus from simply building the most accurate prediction models to evaluating the consequences of providing specific model recommendations to human decision-makers (Fernández-Loría and Provost, 2022b, ; Vodrahalli et al., , 2022). This introduces new challenges for model developers, who need to ensure that the output of an ADM system can be effectively used by a human decision-maker.
@@
For the interaction between human decision-makers and ML models to work, model predictions and its functioning must be comprehensible for human users (Yeomans et al., , 2019; Nourani et al., , 2019; Amarasinghe et al., , 2023). For instance, Lebovitz et al., (2022) show how opaque ML models make it more difficult for medical professional to effectively use them for diagnosis. Many methods have been proposed to make ML models interpretable and explainable, with comprehensive overviews available in Belle and Papantonis, (2021); Molnar, (2022); Doshi-Velez and Kim, (2017); Murdoch et al., (2019); Rudin et al., (2022). One of the reasons why these approaches vary widely is because they have very different conceptions of what constitutes an understandable explanation of the output of a model. Amarasinghe et al., (2023) establish an initial taxonomy linking public policy use cases with existing explainable ML approaches. Moving forward, they stress the need to rigorously evaluate explainable ML methods in real-world problem contexts to ensure their effectiveness in achieving real policy goals and in aiding domain experts. Research from fields such as psychology, cognitive sciences and philosophy may help in the task of creating explanations that are helpful to human users (Miller, , 2019). This requires careful investigation of various challenges, such as identifying situations where model explanations may be harmful due to information overload (Poursabzi-Sangdeh et al., , 2021), or when users may take advantage of increased transparency to exploit a system (Molnar, , 2022). Similarly, misleading explanations can be used to manipulate users and unjustifiably increase trust in a system (Lakkaraju and Bastani, , 2020).
@@
3 Expanding the ADM Toolkit: Choosing the Target Estimand
@@
In Section 2, we outlined several challenges that could threaten the intended functioning of an ADM system using supervised ML models to inform public sector decision-making. These challenges highlight several limitations of solely relying on a traditional supervised ML framework in model design (Wang et al., , 2023). In response to these limitations, there have been calls to move beyond purely predictive modeling towards ML methodology that more directly centers around the goal of decision-making (Hüllermeier, , 2021). This involves a shift in perspective from solely focusing on achieving accurate predictions to a more holistic modus operandi centered on selecting a modeling approach that can best inform the decision-making for a given policy goal and application context.
To illustrate this shift, we will discuss three distinct modeling frameworks, starting with standard risk prediction, which is commonly used in ADM systems, and then move to explore two additional causal modeling frameworks. Standard (risk) prediction (Section 3.1) focuses on estimating outcomes based on historical data without explicitly considering causality. Counterfactual modeling (Section 3.2) extends this approach by estimating causal outcomes of different hypothetical decisions, directly addressing issues such as the influence of past decision-making on the available outcome data. Lastly, policy learning (Section 3.3) aims to directly learn decision policies that maximize a predefined overarching utility, offering a practical approach to optimize a decision policy within the constraints of real-world scenarios (Section 2.4). Figure 2 visually compares how well counterfactual prediction and policy learning address the challenges we have discussed in the context of standard prediction.
Figure 2: Comparing ML Frameworks for Public Sector Decision-Making. Green lines indicate where a causal ML framework is potentially better at addressing a challenge, red lines highlight additional difficulties, and gray lines represent the baseline difficulty of using standard predictive modeling.
@@
First, our goal is to examine the implicit and explicit assumptions underlying each modeling framework. This involves addressing two questions: 1) whether the target estimand in each approach is sufficiently linked to the decision-making process, meaning it would genuinely aid in making informed decisions. Understanding these connections is complex, as the guiding principles of an ADM system can be ambiguous, even when specific goals are in place. For example, public employment agencies often seek to allocate resources to job seekers at higher risk of long-term unemployment. However, this objective might stem from either a belief in the effectiveness of early interventions for high-risk job seekers or the notion that high-risk job seekers inherently deserve more support (Desiere and Struyven, , 2021). Such ambiguity can present difficulties, complicating the choice of the appropriate target estimand, as a need-based distribution necessitates different modeling considerations than an approach focused on the most efficient allocation of interventions. 2) whether estimation is feasible, and what external assumptions are necessary to ensure the accuracy of such estimates. We will discuss these questions for each framework, allowing decision-makers to assess the validity of each approach for their application context.
@@
We will then discuss how the (remaining) challenges outlined in Section 2 can be addressed within each framework. We will present methodological advancements specific to each modeling approach that can help overcome the discussed challenges. While some challenges may be common across all frameworks, others might be more pronounced in specific modeling approaches. Specifically, we focus on distribution shifts to ensure robustness across deployment environments, uncertainty quantification as a key building block for generating trustworthy predictions for humans, and multi-objective optimization to manage tradeoffs between competing objectives.
While we highlight three central modeling approaches, it is important to note that we do not address modeling approaches for every potential decision-making setting. Scenarios involving continuous interventions, interventions across time, sequential data, or interventions that simultaneously target multiple outcomes require other specialized approaches, which are beyond the scope of this paper (Hüllermeier, , 2021; Lin et al., , 2021; Acharki et al., , 2023; Van Geloven et al., , 2020).
@@ 3.1.1 Distribution Shifts, Selection Bias and Label Bias
Most supervised learning models assume that training and deployment data follow the same distribution. However, in many real-world scenarios we may encounter a distribution shift between the training and deployment environment, necessitating the development of reliable models capable of handling and mitigating such differences (Duchi and Namkoong, , 2021). Training models that remain valid under distribution shift often requires assumptions about the expected type of shift (David et al., , 2010). As outlined in Section 2.1, we will predominantly focus on covariate, label and concept shifts (Moreno-Torres et al., , 2012; Quinonero-Candela et al., , 2008). Additionally, we will explore label bias as a type of distribution shift introduced through the use of proxy variables, and consider shifts in time caused by non-stationary environments. We outline central research streams below – for more in-depth introductions to transfer learning, domain adaptation and out-of-distribution generalization approaches, see Kouw and Loog, (2018) and Zhou et al., (2022).
The survey research literature is an invaluable resource for understanding and systematizing different error sources in the data collection process that may surface as distribution shifts downstream. With their inherent focus on valid population inference, concepts such as under-coverage (relevant subpopulations cannot be reached with a data collection schema) and non-response (potentially selective non-participation of relevant subpopulations) extend beyond the traditional survey setting and can help to systematize deficits in the training data in relation to the target population. Error frameworks such as the Total Survey Error (Groves and Lyberg, , 2010; Biemer, , 2010) and its extensions (Sen et al., , 2021) have been proposed to systematically trace errors along the data collection and processing pipeline that can accumulate in misrepresentation issues. Related work proposes strategies for improving inference from data that do not adequately represent the target population of interest (Cornesse et al., , 2020; Yang and Kim, , 2020). In this context, pseudo-weighting approaches (Elliott and Valliant, , 2017; Valliant and Dever, , 2011) are employed to match the potentially biased source data to some known reference distribution, which resembles methodology from the domain adaptation literature (see below) and similarly connects to concepts in causal inference (Mercer et al., , 2017). As a recent example of cross-disciplinary work in this context, Kim et al., (2022) draw on the multicalibration framework from algorithmic fairness (Hebert-Johnson et al., , 2018) to learn prediction functions that are universally adaptable to unknown deployment shifts.
@@
However, weighting methods struggle in settings with limited and complex source data, frequently resulting in high variance estimates, and depend on data being available from the target domain of interest (Fang et al., , 2020; Kouw and Loog, , 2018; Liu and Ziebart, , 2014). Distributionally robust methods offer an alternative approach by providing worst-case guarantees. They often involve minimax estimation, which seek to minimize the loss under the least favorable distribution shift (Kouw and Loog, , 2018; Subbaswamy et al., , 2022; Duchi and Namkoong, , 2021; Wen et al., , 2014).
@@
Various approaches have been devised to mitigate label noise and measurement error. For instance, Natarajan et al., (2013) propose an unbiased risk minimization strategy for handling class-conditional $p(\tilde{Y}|Y)$ noise. While such a simplified model of a proxy may be applicable in some settings, practitioners will likely encounter more complex scenarios (Chen et al., , 2021), such as the measurement error depending on sensitive covariates (Wang et al., , 2021; Obermeyer et al., , 2019). In some situations, there may be the option to access multiple proxies of the true target of interest. For example, Boeschoten et al., (2021) utilize a structural equation model to characterize the relationship between multiple proxies and the unobserved outcome to ensure fair predictions. There has also been research into how label bias interacts with other distribution shifts. For example, Dai and Brown, (2020) propose a joint framework for addressing label bias and label shift, while Yu et al., (2020) investigate the interaction between class-conditional noise and generalized target shifts.
Distribution shifts tend to occur gradually over time (Webb et al., , 2016), often triggered by the deployment of the model itself. Addressing such feedback loops and ongoing distribution shifts poses a significant challenge, likely requiring future research into the temporal dynamics of ML-informed decision-making (Pagan et al., , 2023). For example, Perdomo et al., (2020) introduce a modeling framework that incorporates the potential impact of predictions on the predicted outcome of interest. These predictions are referred to as performative, effectively leading to distribution shifts by altering the target distribution in the deployment environment over time. They develop the notion of performative optimality, ensuring that a decision rule minimizes the expected loss with regard to the future target distribution it induces. The specific choice of the loss function can align with different objectives. For instance, one may opt to optimize for a target distribution with mostly favorable outcomes instead of solely focusing on accurate predictions (Kim and Perdomo, , 2023).